OpenStudy (anonymous):

A farmer examines a sample of 25 cartons of eggs and finds that 3 contain cracked eggs. What is the best prediction of the number of cartons with cracked eggs in a delivery with 500 cartons.

OpenStudy (anonymous):

3 cracked eggs per 25 eggs 3/25 so divide 500/25 you will get 20 multiply 3/25 both top and bottom by 20 you will get 60/500 or 3/5

OpenStudy (anonymous):

This means 60 eggs out of the 500 will be cracked:) hope this helps!

OpenStudy (anonymous):

that is a good explanation for a quick computation you can write \[\frac{3}{25}=\frac{x}{500}\iff x=\frac{500\times 3}{25}\] and compute that number